The rate constant for the reaction H^+ + OH^- righ
Answer #1

Given reaction H+ + OH- -----------> H2O

rate constant k = 1.3 x 1011 dm3 mol-1 s-1 =  1.3 x 1011 L mol-1 s-1 ( 1dm3 = 1L)

The given reaction is 2nd order reaction because units for rate constant of 2nd order reaction =  L mol-1 s-1

   Let us assume a 2nd order reaction, A + B ------------> P

If initial concentrations of A and B are same, expression for half-life is

half-life = 1/(k[Ao]) where Ao = initial concentration of A

(a) [H+] = [OH-] = 10-1 M = 10-1 mol L-1

[Ao] = 10-1 mol L-1

Hence, half-life = 1/(k[Ao])

= 1/ (1.3 x 1011 L mol-1 s-1 x 10-1 mol L-1)

= 0.77 x 10-10 s

Therefore, half-life = 0.77 x 10-10 s

b) [H+] = [OH-] = 10-1 M = 10-1 mol L-1

[Ao] = 10-1 mol L-1

Hence, half-life = 1/(k[Ao])

= 1/ (1.3 x 1011 L mol-1 s-1 x 10-1 mol L-1)

= 0.77 x 10-10 s

Therefore, half-life = 0.77 x 10-10 s